What is the slope of a line that is perpendicular to a line that has no slope? Explain the answer.
Answer by drk(1908)

If a line has no slope then it can be expressed as 1/0 or in general as a/0, where a not = 0.
A perpendicular slope is the negative reciprocal.
So, 1/0 turns into -0/1 = 0,
or in general a/0 turns into -0/a = 0.